H3PO4 is a tribasic acid and one of its salts is NaH2PO4.What volume of 1 M NaOH should be added to 12 g of NaH2PO4(MOl.Wt. 120) to exactly convert into Na3PO4

Grade:12th pass

1 Answers

Aman Bansal
10 years ago

Dea Vempally,

Balanced equation:
NaH2PO4 + 2NaOH → Na3PO4 + 2H2O
1mol NaH2PO4 reacts with 2mol NaOH
Molar mass NaH2PO4 = 120.0 g/mol
6g NaH2PO4 = 6/120 = 0.05mol NaH2PO4
This will require 0.05*2 = 0.10 mol NaOH
3mol NaOH = 1000mL 
0.10mol = 0.10/3*1000 = 33.3mL NaOH required.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert